system updates, 2011-02-11
Tom Krajci informed us that W30 had a filter wheel malfunction of 2011-02-10UT. It was fixed on the 11th, but we need to reprocess 110210 and 110208 to ensure that those images were not compromised. Please hold off on submitting data from W30 for those two nights until we post the results. The weather is currently clear, so there will be new images for the next few nights, but since we are entering a weekend, you may not see new images posted until Monday.
John Gross indicates that SRO50 had a dome problem last night (110211). Some of the images may be influenced; we need to look at them and make a report. He will be going down to Sonoita tomorrow to look at the situation.
It has been cloudy for the past week or so at Astrokolkhoz, so you will not see any images between 110131 and 110207 inclusive.